# GECAM Observations of the Galactic Magnetar SGR J1935+2154 during the 2021 and 2022 Burst Active Episodes. I. Burst Catalog

### v5

@@ 3.4 Burst Waiting Time
Fig. 10 shows the distribution of waiting time measured by GECAM-B/C. These distributions could be well fitted with lognormal functions, resulting in a mean value of seconds for GECAM-B and a mean value of seconds for GECAM-C. Compared to a lognormal distribution, the Kolmogorov-Smirnov tests yield a statistic of 0.104 and a p-value of 0.672 for GECAM-B, and a statistic of 0.112 and a p-value of 0.929 for GECAM-C. This lognormal distribution behavior is similar to that observed in previous studies (Cai et al., 2022a; Xie et al., 2024). Since the relatively short continuous observation time intervals of GECAM-C for observing SGR J1935+2154 (only dozens of minutes) compared to GECAM-B, it is difficult to measure the longer waiting times between more successive bursts. Therefore, the distribution’s mean value of GECAM-C is smaller than that of GECAM-B.
3.5 Burst Hardness Ratio
@@ 4 Summary and Discussions
The burst activity is evaluated by estimating the daily burst rate data, as illustrated in Fig. 5. We employ the Lomb-Scargle method to investigate potential periodic behavior, which results in the most likely period of days (refer to Fig. 6). Therefore, the burst history over these two years could be divided into four active episodes (see Fig. 5). The period of days with 80% duty cycle is well consistent with previous study (Xie et al., 2022, evaluated from July 2014 to January 2022). Unlike other periods derived in previous work ( days evaluated from 2014 to 2020 or days evaluated from July 2014 to October 2021 in Grossan, 2021; Zou et al., 2021, respectively), a shorter period is found in this study, possibly indicating an increased activity of X-ray bursts from SGR J1935+2154 after the year of 2021.
@@
We perform extensive studies on the statistical characteristics of these X-ray bursts, including the burst duration, waiting time, and hardness ratio. The duration follows a lognormal distribution for GECAM-B/C, respectively (see Fig. 9 and Table 5). Due to the varying effective detectable energy ranges of GECAM-B and GECAM-C, some bursts detected by both satellites may show different durations (see Fig. 4 for examples). In such cases, the duration measured by GECAM-C is more informative across the entire energy range when both satellites detect the same burst. The burst duration of SGR J1935+2154 is around 100 ms, see Table 6 for comparison among different burst episodes. The waiting time also conforms to a lognormal function with mean values of seconds and seconds for GECAM-B/C, respectively (refer to Fig. 10). Because of the relatively short continuous observation time intervals for GECAM-C observing SGR J1935+2154 compared to GECAM-B, the distribution’s mean value of GECAM-C is smaller than that of GECAM-B. Such lognormal distribution behavior is similar to that observed in previous studies (Cai et al., 2022a; Xie et al., 2024). The hardness ratio of X-ray bursts tends to become softer over the course of the observation time spanning these two years (see Fig. 11).
These statistical results indicate that the X-ray bursts from SGR J1935+2154 show increased activity, as observed by Fermi/GBM (e.g., Rehan & Ibrahim, 2024). The magnetar bursts also exhibit a softer spectrum over the two-year period, as presented by the temporal evolution of the hardness ratio and a detailed spectral analysis will follow. Similar to FRB 200428, which occurred during the most active episode (April-May 2020) preceding 2020, the radio burst FRB 20221014 are detected in the final and most active episode (October-December 2022). In view of their millisecond duration and high energy releases, FRBs are widely suggested to originate from the violent activities of compact objects, in particular magnetars (Popov & Postnov, 2010; Kulkarni et al., 2014; Katz, 2016; Connor et al., 2016; Cordes & Wasserman, 2016; Lyutikov, 2017). Leading to such burst activities may correspond to some sudden changes in the magnetic configuration of the magnetars (e.g., starquakes, magnetic field reconnections Thompson & Duncan, 1995; Thompson & Blaes, 1998; Jones, 2003; Levin & Lyutikov, 2012). In that case, FRB emission may result from such burst interactions (e.g., the collision between different Alfven waves or different explosion outflows) in a global activity of a magnetar and then a detected FRB be associated with an X-ray burst (Yang & Zhang, 2021; Xie et al., 2024).

### v6

@@ III.4 Burst Waiting Time
Fig. 10 shows the distribution of waiting time measured by GECAM-B/C. These distributions could be well fitted with lognormal functions, resulting in a mean value of seconds for GECAM-B and a mean value of seconds for GECAM-C. Compared to a lognormal distribution, the Kolmogorov-Smirnov tests yield a statistic of 0.104 and a p-value of 0.672 for GECAM-B, and a statistic of 0.112 and a p-value of 0.929 for GECAM-C. This lognormal distribution behavior is similar to that observed in previous studies (Cai2022apjs_A; Xie2024ApJ). Since the relatively short continuous observation time intervals of GECAM-C for observing SGR J1935+2154 (only dozens of minutes) compared to GECAM-B, it is difficult to measure the longer waiting times between more successive bursts. Therefore, the distribution’s mean value of GECAM-C is smaller than that of GECAM-B.
III.5 Burst Hardness Ratio
@@ IV Summary and Discussions
The burst activity is evaluated by estimating the daily burst rate data, as illustrated in Fig. 5. We employ the Lomb-Scargle method to investigate potential periodic behavior, which results in the most likely period of days (refer to Fig. 6). Therefore, the burst history over these two years could be divided into four active episodes (see Fig. 5). The period of days with 80% duty cycle is well consistent with previous study (Xie2022mnras, evaluated from July 2014 to January 2022). Unlike other periods derived in previous work ( days evaluated from 2014 to 2020 or days evaluated from July 2014 to October 2021 in Grossan2021PASP; Zou2021apjl, respectively), a shorter period is found in this study, possibly indicating an increased activity of X-ray bursts from SGR J1935+2154 after the year of 2021.
@@
We perform extensive studies on the statistical characteristics of these X-ray bursts, including the burst duration, waiting time, and hardness ratio. The duration follows a lognormal distribution for GECAM-B/C, respectively (see Fig. 9 and Table 5). Due to the varying effective detectable energy ranges of GECAM-B and GECAM-C, some bursts detected by both satellites may show different durations (see Fig. 4 for examples). In such cases, the duration measured by GECAM-C is more informative across the entire energy range when both satellites detect the same burst. The burst duration of SGR J1935+2154 is around 100 ms, see Table 6 for comparison among different burst episodes. The waiting time also conforms to a lognormal function with mean values of seconds and seconds for GECAM-B/C, respectively (refer to Fig. 10). Because of the relatively short continuous observation time intervals for GECAM-C observing SGR J1935+2154 compared to GECAM-B, the distribution’s mean value of GECAM-C is smaller than that of GECAM-B. Such lognormal distribution behavior is similar to that observed in previous studies (Cai2022apjs_A; Xie2024ApJ). The hardness ratio of X-ray bursts tends to become softer over the course of the observation time spanning these two years (see Fig. 11).
These statistical results indicate that the X-ray bursts from SGR J1935+2154 show increased activity, as observed by Fermi/GBM (e.g., Rehan2024ApJ). The magnetar bursts also exhibit a softer spectrum over the two-year period, as presented by the temporal evolution of the hardness ratio and a detailed spectral analysis will follow. Similar to FRB 200428, which occurred during the most active episode (April-May 2020) preceding 2020, the radio burst FRB 20221014 is detected in the final and most active episode (October-December 2022). In view of their millisecond duration and high energy releases, FRBs are widely suggested to originate from the violent activities of compact objects, in particular magnetars (Popov2010; Kulkarni2014; Katz2016; Connor2016; Cordes2016; Lyutikov2017). The mechanisms leading to such burst activities may correspond to some sudden changes in the magnetic configuration of the magnetars (e.g., starquakes, magnetic field reconnections Thompson1995MNRAS; Thompson1998PhRvD; Jones2003ApJ; Levin2012MNRAS). In that case, FRB emission may result from such burst interactions (e.g., the collision between different Alfven waves or different explosion outflows) in a global activity of a magnetar and then a detected FRB be associated with an X-ray burst (Yang2021ApJ; Xie2024ApJ).